A local salesman receives a base salary of $800 monthly. He also receives a commission of 9% on all sales over $1200. How much would he have to sell in a month if he needed to have a monthly income of $1800?
Answer by mananth(16946)

salary 800
let sales =x

commission = 9% on sales above 1200
=0.09*(x-1200)

Then I = 800+0.09*(x-1200)
1800=800+0.09x-108
1000-800+108=0.09x
308=0.09x
308/0.09=x
x=3422
He will have to make sales of 3422+800=$4222.22